OpenStudy (anonymous):

limx->0 (x/abs(x)) . sinx

OpenStudy (anonymous):

seems unlikely

OpenStudy (anonymous):

\[\frac{x}{|x|}\] is 1 if \(x>0\) and \(-1\) if \(x<0\)

OpenStudy (anonymous):

Why is that? Wouldn't Sin(0) = 0 ?

OpenStudy (anonymous):

oooh my bad (as we say) i didn't see the \(\sin(x)\) there!!

OpenStudy (anonymous):

it is zero for sure, but there is something to say, because your expression is not defined at \(x=0\)

OpenStudy (anonymous):

you have to take a right and a left sided limit \[\lim_{x\to 0^+}\sin(x)=0\] and \[\lim_{x\to 0^-}\sin(x)=0\]

OpenStudy (anonymous):

you have to take them both because your expression is \(\sin(x)\) if \(x>0\) and it is \(-\sin(x)\) if \(x<0\)

OpenStudy (anonymous):

so ... The answer should be like this? Right? limx→0+sin(x)=0 and limx→0−sin(x)=0 @@satellite73

OpenStudy (anonymous):

we can use this relation \[|\sin(x)|\leq x\] and then multiply by \[|\frac{x}{|x|}|\] we will obtain \[|\frac{x}{|x|}\sin(x)|\leq x\] and then the limit as \[x\to 0\] the limit is zero

OpenStudy (anonymous):

@Mega96 yes